SpaceRef

SpaceRef


ISS On-Orbit Status 12 Mar 2003

Status Report From: NASA HQ
Posted: Wednesday, March 12, 2003

image All ISS systems continue to function nominally, except as noted previously or below.

FE-1 Nikolai Budarin began the day with his routine inspection, including watering if required, of the Russian BIO-5 Rasteniya-2/Lada-2 plant growth experiment which investigates growth and development of acacia-leaf type pea seedlings under spaceflight conditions.

Afterwards, Budarin worked in the DC-1 "Pirs" docking module, removing and replacing two auxiliary digital commutator (TA253B) electronics boxes, a subsystem of the ASS message acquisition system, with each device requiring about 1.5 hrs crewtime.

FE-2/SO Don Pettit continued his unflagging support of the ongoing MSG (Microgravity Science Glovebox) troubleshooting.  [Once again, early analysis showed no off-nominal current draw on the ESEM3 (exchangeable standard electronic module 3) after yesterday's re-installation of the P104 and P110 connectors.  Today, Pettit was to reconnect E-Box connector P112, powering Work Volume illumination units, and, if successful, connector P113, for an E-Box daughter board.  Besides the sensor data downlink via Ku-band, numerous onboard LEDs (light-emitting diodes) are now available to indicate function status to Pettit.]

Later, Pettit and CDR Ken Bowersox completed another DOUG (dynamic operations ubiquitous graphics) computer session in preparation for tomorrow's SSRMS/Robotics operations.  [During last week’s survey of the S1 RBVM (radiator beam valve module) some additional thermal protection "booties" were noted as possibly being open.  Tomorrow's procedure will "walk" the SSRMS to MBS PDGF1 (Mobile Base System/Power & Data Grapple Fixture #1) and maneuver the arm to two viewing positions to inspect these booties.  From the second survey position a general survey of the RBVMs will then be performed, using the Tip Elbow camera, followed by two checkouts of the Tip LEE (latching end effector).  Operations will end with the SSRMS being walked back to the Lab PDGF and maneuvered into position to view the upcoming stage EVA #2.  All OCASs (operator commanded auto sequences) will be performed in vernier rates due to tight clearances with the UHF antenna, Lab camera and S1 radiator, of less than 5 feet during parts of the maneuvers.]

Nikolai Budarin worked about an hour in the Service Module (SM), using the V/K-1 video camera and GNShK low-noise headset microphone to test the newly reconfigured audio recording of the LIV television system for incorporation of this mode into crew procedures.

Budarin later terminated the LSO "Sprite"-seeking experiment, dismantled the "Molniya-SM" payload at SM window #3 and deactivated the French EGE laptop.  All equipment was then placed in stowage.

Nikolai also completed the daily routine maintenance of the SM's SOZh life support system, including ASU toilet facilities, while Bowersox prepared the daily "delta" file of the IMS (inventory management system) database and conducted the regular Lap payload status checkup.

Bowersox downloaded data files from the CMS HRM (crew medical systems/heart rate monitor) to the MEC (medical equipment computer) for subsequent downlink via OCA comm.

All crewmembers performed their daily physical exercise program on RED (resistive exercise device), TVIS (treadmill with vibration isolation and stabilization), CEVIS (cycle ergometer with vibration isolation) and VELO with load trainer.  [With the Stage EVA delayed until early April (see yesterday's Status), use of CEVIS (with failed control display) is OK again for regular crew exercise, which today replaced the originally scheduled pre-EVA PFEs (periodic fitness evaluations).]

MCC-H began the planned uploading of the new R3 GNC (guidance, navigation & control) "stale RGA data" software patch, in conjunction with today's handover of ISS attitude control (AC) to the Russian segment (RS) at 2:10pm EST for the first Progress thruster test reboost tonight at 5:58pm, and the return of AC from RS to USOS (U.S. segment) thereafter.  [The reboost, of 600 seconds duration (primary mode) is planned to produce a delta-V of 1.39 m/sec.  The burn could be changed to backup mode (800 sec), since primary mode shows a conjunction with a piece of space debris some nine hours later, with 18 km miss distance (a spent Soviet SL-12 rocket body/auxiliary motor, Object #13608).  The reboost is controlled by an automatic sequencer, preloaded with a command sequence ("cyclogram") uplinked by MCC-M earlier today during a Russian comm pass on Daily Orbit 11.]

The GNC software upload also includes the new PPL (pre-positioned load) file with larger rate divergence threshold values as an interim solution to the RGA (rate gyro assembly) dilemma issue observed last week (see 3/8 Status Report).

For the popular CEO (crew earth observations program), the crew has requested target data for cities at night, as well as for opportunities to photograph the Nasca Lines, Mayan and Aztec pyramids, Great Wall of China, Angel Falls Venezuela and other challenging man-made and natural features.  The requests are being incorporated in the daily targeting.  Today's CEO targets were Cape Town, South Africa(crew was to document the entire city expanse), Atlanta, Georgia (a nice nadir pass to capture city borders), El Paso, Texas (crew was to look slightly right of track for a good clear view of El Paso), Patagonian Glaciers (clear weather for great views.  Shooting detailed glacier features), and Maro (image slightly right of track to get reef structure). CEO images can be viewed at the websites http://eol.jsc.nasa.gov and http://earthobservatory.nasa.gov

U.S. and Russian Segment Status(as of 1:55pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (24 Amp mode).  Vozdukh CO2 scrubber is On (Manual mode 5).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode; bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) -- 753; temperature (deg C) -- 26.9; ppO2 (mmHg) -- 184.1; ppCO2 (mmHg) -- 2.5 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) -- 751; temperature (deg C) -- 19.8.
  • FGB Cabin:  Pressure (mmHg) -- 752; temperature (deg C) -- 20.3.
  • Node:  Pressure (mmHg) -- 743.25; temperature (deg C) -- 22.7 (shell); ppO2 (mmHg) -- 171.0; ppCO2 (mmHg) -- 4.4.
  • U.S. Lab:  Pressure (mmHg) -- 745.52; temperature (deg C) -- 22.2; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 745.52; temperature (deg C) -- 20.8; shell heater temp (deg C) -- 22.4, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • PMA-1:  Shell heater temp (deg C) -- 22.9
  • PMA-2:  Shell heater temp (deg C) -- 20.9

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available [SM(820) + FGB(2746) + Progress(576)] -- 4142 kg (9132 lb)as of 3/6/03.  (Capability: SM -- 860 kg; FGB -- 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and 4B both in dual-angle directed position (eclipse drag reduction configuration, "night glider").
  • SM batteries:  All batteries (8) are in "Partial Charge" mode.
  • FGB batteries:  Batteries #2 is off line; all other batteries (5) are in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is On (primary); PL-1 MDM is Off
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.  Lane 1 is out of the set (as of 11/14/02).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.  Lane 1 is out of the set (as of 3/2/03).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- Russian
  • Angular rate source -- Russian.

Flight Attitude:

  • LVLH TEA (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.1 deg, roll: 0 deg]), with CMG/ TA (thruster assist) Momentum Management).
  • Solar Beta Angle:  20.2 deg (magnitude increasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally (VTR1 is operable again).
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched at WS4, with KA power. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:21am EST [= epoch]):

  • Mean altitude -- 388.7 km
  • Apogee -- 396.7 km
  • Perigee -- 380.7 km
  • Period -- 92.33 min.
  • Inclination (to Equator) --  51.63 deg
  • Eccentricity -- 0.0011847
  • Orbits per 24-hr. day -- 15.59
  • Solar Beta Angle:  20.2 deg (magnitude increasing)
  • Mean altitude loss in last 24 hours -- 150m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  -- 24592
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •   http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.